A survey of 350 residents of a certain city found that 82 used public transportation daily. You wish to find a 95% confidence interval for the true proportion of residents who use public transportation daily. What is the lower limit of the interval?
Answer #1

p = 82 / 350 = 0.234

Z for 95% confidence interval = Z0.025 = 1.96

Lower limit = p - Z0.025 * sqrt(p * (1 - p) / n)

                  = 0.234 - 1.96 * sqrt(0.234 * (1 - 0.234) / 350)

                  = 0.190